OpenStudy (anonymous):

the max value of f(x)= 2x^3-9x^2+12x-1 on [-1,2] is

OpenStudy (fibonaccichick666):

any ideas on how to begin?

OpenStudy (anonymous):

mm, take the first deriv and set equal to 0? i know that's what you do to find max value... but, idk i can't seem to get the answer correct

OpenStudy (fibonaccichick666):

ok that's a good start

OpenStudy (fibonaccichick666):

setting = 0 will tell you the critical points

OpenStudy (fibonaccichick666):

did you remember to check the end pts?

OpenStudy (anonymous):

which are the end points? >.< lol

OpenStudy (anonymous):

\[f'\left( x \right)=6x^2-18x+12=0~gives~x^2-3x+2=0,x^2-2x-x+2=0\] \[x \left( x-2 \right)-1\left( x-2 \right)=0,\left( x-2 \right)\left( x-1 \right)=0,x=1,2\] check at x=-1,1,2

OpenStudy (fibonaccichick666):

what is your domain?

OpenStudy (anonymous):

where is the @surjithayer -1 from?

OpenStudy (anonymous):

and how do you check them? :0

OpenStudy (anonymous):

jst substitute them back into the equation??

OpenStudy (fibonaccichick666):

test them, I suggest finding the critical points, then seeeing if the first deriv is neg or pos

OpenStudy (anonymous):

you are given the interval[-1,2] i found two values x=1,x=2

OpenStudy (anonymous):

okay, got that! but how do you test the x values? it's a non calculator section...

OpenStudy (fibonaccichick666):

you want to look at the first derivative, what does it represent?

OpenStudy (anonymous):

\[f \left( x \right)=2x^3-9x^2+12x-1,find f \left( -1 \right),f \left( 1 \right),f \left( 2 \right)\]

OpenStudy (anonymous):

you got that eqn by taking the intergral ..?

OpenStudy (fibonaccichick666):

@SoccerGirl13 Let's restart, I think you may be getting a little confused. What was your given?

OpenStudy (anonymous):

f(x)= 2x^3-9x^2+12x-1

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (fibonaccichick666):

alright, and what are we evaluating it over?

OpenStudy (anonymous):

[-1,2]

OpenStudy (fibonaccichick666):

good, so the thing to remember is that you always have to check your end points and any critical points on your domain, do you follow?

OpenStudy (anonymous):

yea :D I get it (: it all just clicked :)

OpenStudy (anonymous):

thanks (:

OpenStudy (fibonaccichick666):

np, also, when you can't graph. Check your first derivative, is it positive? that means your graph has a positive slope. Which means increasing, if it is increasing on your domain, you know that the end pt must be the local max